OpenStudy (anonymous):

The Question says to use a non-calculator method to calculate the following binomial coefficients. I will post the first question now. I do not know any non-calculator method and cant find one anywhere so please help!

OpenStudy (anonymous):

(vi) \[\left(\begin{matrix}12 \\ 9\end{matrix}\right)\]

OpenStudy (misty1212):

cancel first, multiply last, you generally don't need a calculator since the binomial coefficients are whole number

OpenStudy (misty1212):

\[\binom{12}{9}=\binom{12}{3}\] is a good start to make your life easier

OpenStudy (misty1212):

then \[\binom{12}{3}=\frac{12\times 11\times 10}{3\times 2}\]

OpenStudy (jango_in_dtown):